OpenStudy (anonymous):

help me solve x=pi/2+k(pi), -3(pi) less than or equal to x and x is also less than or equal to 3(pi). k is any integer list all possible values for x that are specified in interval.

OpenStudy (anonymous):

Let n=pi x=n/2+k*n, -3n<=x, x<=3n So, -3n<=n/2+k*n --> (-3n-n/2)/n <= k And, n/2+k*n<=3n --> k<=(3n-n/2)/n Therefore (-3n-n/2)/n <= k <=(3n-n/2)/n Those are your bounds for k, substitute each into x=n/2+k*n and you'll have your bounds for x.

OpenStudy (anonymous):

ok thank you